Circular No. 6061

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

γ-RAY FLARE FROM DIRECTION OF NRAO 190 T. A. McGlynn and J. R. Mattox, Compton Observatory Science Sup-

port Center, Goddard Space Flight Center, NASA; W. T. Vestrand, University of New Hampshire; and the Compton Observatory EGRET Team (B. L. Dingus, D. L. Bertsch, J. A. Esposito, C. E. Fichtel, R. C. Hartman, S. D. Hunter, C. von Montigny, R. Mukherjee, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; E. Schneid, Grumman Aerospace Corporation; and G. Kanbach and H. A. Mayer-Hasselwander, Max-Planck-Institut für Extraterrestrische Physik) communicate: “We report the detection of a γ-ray source, having a position consistent with the quasar NRAO 190 (Pauliny-Toth et al. 1966, Ap.J. Supp. 13), in the energy range above 100 MeV during Aug. 9–20. The maximum-likelihood position from the EGRET quicklook data is l = 197^{o}.0, b = −29^{o}.5 (α = 4^{h}38^{m}.7, δ = −0^{o}41′, equinox 2000.0), with a 95-percent error cone having a radius of at least 1^{o}. The proposed identification is 1^{o} from the γ-ray position and is the only strong radio-loud active galactic nucleus within the 95-percent confidence contour. EGRET will continue to view this region of the sky until Aug. 29. The identification is consistent with the typical EGRET detection of strong, flat-spectrum, optically-violently variable QSOs. The outburst has a flux at least four times the 95-percent upper limit of 8 × 10^{−8} count cm^{−2} s^{−1} (E > 100 MeV) obtained during the EGRET all-sky survey. We urge observations at other wavelengths to attempt to determine the characteristics of the outburst.”

NOVA OPHIUCHI 1994 R. J. Ivison and D. Ward-Thompson, Royal Observatory, Edinburgh;

and H. M. Lloyd, Liverpool John Moores University, on behalf of the James Clerk Maxwell Telescope nova-monitoring group, report the detection of Nova Oph 1994: “The flux density at 1.1 mm was 29 ± 9 mJy on June 30.4 UT. On Aug. 7.4, the flux density at 0.8 mm was 110 ± 22 mJy, while that at 1.1 mm was below 61 mJy (3σ), suggesting that the ejecta are still optically thick.”

1994 August 26 Daniel W. E. Green